# Spin-Seebeck Signatures of Spin Chirality in Kagome Antiferromagnets

### v1

@@ Abstract
Non-collinear antiferromagnets (NCAFMs) are appealing for antiferromagnetic spintronics, as they combine the advantages of collinear antiferromagnets with novel emergent phenomena stemming from their complex spin structures. These phenomena are often associated with the intrinsic spin chirality, which characterizes the handedness of the ground-state spin configuration. Here, we investigate a kagome NCAFM interfaced with a normal metal and demonstrate that the ground-state vector spin chirality can be probed through measurements of the spin Seebeck effect (SSE). Starting from a microscopic spin Hamiltonian, we derive the corresponding bosonic Bogoliubov-de Gennes Hamiltonians for the two chiral configurations. Using linear response theory, we obtain a general expression for the spin current thermally pumped into the normal metal by the SSE. We show that a sizable in-plane spin current emerges exclusively in the negative-chiral state, providing a direct signature for real-time detection of chirality switching in kagome NCAFMs. In addition, we find a field-dependent out-of-plane spin current whose magnitude differs between the two chiralities by about 4%, reflecting their distinct magnon band structures.
@@ I Introduction
Chiral magnetic orders have garnered significant attention in the study of non-collinear antiferromagnets (NCAFMs) — antiferromagnetic systems in which the spin sublattices are ordered non-collinearly [5]. The complexity and chirality of their magnetic states are believed to give rise to unconventional magnetic, electronic, and optical phenomena, even in the absence of a net magnetization [6, 7, 8, 9, 10, 11, 12, 13]. A particularly attractive subclass of NCAFMs for exploring chirality-driven spin phenomena is the kagome antiferromagnets (AFMs). Kagome AFMs are two-dimensional (2D) spin systems composed of corner-sharing triangles and are considered among the most geometrically frustrated 2D spin lattices [14] (see Fig. 1). Important examples of materials that can be characterized as kagome AFMs are ultrathin films of the Weyl antiferromagnets [15] $\text{Mn}_{3}\text{X}$ (X = Ge, Sn, Ga) and iron jarosites [14], which consist of stacked kagome layers. In equilibrium, the three sublattice spins ${\boldsymbol{S}}_{1}$, ${\boldsymbol{S}}_{2}$, and ${\boldsymbol{S}}_{3}$ on each triangle are oriented at 120° to one another. The handedness of this spin configuration is described by the vector spin chirality ${\boldsymbol{K}}_{\mathrm{v}}={\boldsymbol{S}}_{2}\times{\boldsymbol{S}}_{1}+{\boldsymbol{S}}_{3}\times{\boldsymbol{S}}_{2}+{\boldsymbol{S}}_{1}\times{\boldsymbol{S}}_{3}$ [14]. For coplanar arrangements, ${\boldsymbol{K}}_{\mathrm{v}}$ points parallel or antiparallel to the z-axis, corresponding to the ground states of $(+)$-chirality and ()(-)-chirality, respectively (see Fig. 1). When subjected to an external magnetic field or an in-plane Dzyaloshinskii-Moriya interaction (DMI), the sublattice spins acquire a slight out-of-plane canting, leading to a non-coplanar spin texture. In this regime, the spin configuration also exhibits a finite scalar spin chirality expressed through the quantity $K_{\mathrm{s}}={\boldsymbol{S}}_{1}\cdot({\boldsymbol{S}}_{3}\times{\boldsymbol{S}}_{2})$ [14].
@@ II Theory
which includes the following contributions: an exchange interaction between nearest-neighbor spins ${\boldsymbol{S}}_{i}$ and ${\boldsymbol{S}}_{j}$ with coupling strength $J>0$, represented by $\mathcal{H}_{\text{E}}=J\sum_{\langle ij\rangle}{\boldsymbol{S}}_{i}\cdot{\boldsymbol{S}}_{j}$, easy-plane ($K_{\perp}>0$) and easy-axis ($K>0$) anisotropy energies, characterized by $\mathcal{H}_{\mathrm{A}}=\sum_{i}\left(K_{\perp}({\boldsymbol{S}}_{i}\cdot\hat{{\boldsymbol{z}}})^{2}-K({\boldsymbol{S}}_{i}\cdot\hat{{\boldsymbol{n}}}_{i})^{2}\right)$, an out-of-plane DMI between nearest neighbors in the form of $\mathcal{H}_{\mathrm{DM}}=\sum_{\langle ij\rangle}D{\boldsymbol{\hat{z}}}\cdot({\boldsymbol{S}}_{i}\times{\boldsymbol{S}}_{j})$, and the Zeeman coupling to an external magnetic field $B$ applied perpendicular to the kagome lattice plane, $\mathcal{H}_{\mathrm{ext}}=-\sum_{i}\gamma B\hat{{\boldsymbol{z}}}\cdot{\boldsymbol{S}}_{i}$, where $\gamma$ is the gyromagnetic ratio. The in-plane easy-axes at the three spin sub-lattice sites are $\hat{{\boldsymbol{n}}}_{1}=[0,1,0]^{T}$, $\hat{{\boldsymbol{n}}}_{2}=[\frac{\sqrt{3}}{2},-\frac{1}{2},0]^{T}$, and $\hat{{\boldsymbol{n}}}_{3}=[-\frac{\sqrt{3}}{2},-\frac{1}{2},0]^{T}$, respectively. In what follows, it will often be convenient to represent a lattice site as $i\equiv(\kappa,\alpha)$, with $\kappa$ denoting the unit-cell index and $\alpha\in\{1,2,3\}$ the sublattice index.
@@
Figure 2: (color online). Magnon energy bands along high-symmetry lines in the first Brillouin zone for a. the $(+)$-chiral ground state and b. the ()(-)-chiral ground state. The inset diagrams highlight the band structures near the $\Gamma$ point along the $x$-axis. We have used the parameter values $2a=5.665$ Å, spin $S=1$, $\gamma=1.76\cdot 10^{11}$ 1/Ts, $J=17.53~\mathrm{meV}/\hbar^{2}$, $196~\mathrm{meV}/\hbar^{2}, K=0K_{\bot}=0$, $D=K/4\sqrt{3}$, and $B=5$ T.
Fig. 2 shows the magnon band structures calculated from Eq. (3) for the ($\pm$)-chiral states. The parameters correspond to $\text{Mn}_{3}\text{Sn}$ with lattice constant $2a=5.665$ Å, spin $S=1$, gyromagnetic ratio $\gamma=1.76\cdot 10^{11}$ 1/Ts, $J=17.53~\mathrm{meV}/\hbar^{2}$, $K_{\bot}=0$, and $K=0.196~\mathrm{meV}/\hbar^{2}$ [33, 48]. For the out-of-plane DMI, we take the threshold value corresponding to the transition between positive and negative chiralities, $D=K/4\sqrt{3}$. Unless otherwise specified, these material parameters are used in all subsequent calculations. Additionally, for the band structures in Fig. 2, an external magnetic field of $B=5$ T is applied.
The main difference between the band structures of the $(\pm)$-chiral states arises near the $\Gamma$ point, as highlighted in the inset diagrams of Fig. 2. In the $(+)$-chiral state, the lowest band exhibits a finite energy gap at ${\boldsymbol{q}}=0$, whereas in the ()(-)-chiral state the lowest band approaches zero as ${\boldsymbol{q}}\to 0$. At ${\boldsymbol{q}}=0$, the band retains only a small finite value determined by the in-plane tilt angle $\alpha$, which vanishes in the strong exchange-coupling limit considered in Ref. [36]. At finite temperature, this results in a larger magnon population in the lowest band of the ()(-)-chiral state compared to the $(+)$-chiral state, thereby leading to an amplification of the SSE, as we demonstrate below.
@@ III The Spin Seebeck Effect
with $J_{\mathrm{sd},i}\equiv\int{\rm d}{\boldsymbol{r}}J_{i}\rho_{i}({\boldsymbol{r}})|\Psi_{{\boldsymbol{R}}_{i}}({\boldsymbol{r}})|^{2}$ and ${\boldsymbol{s}}_{i}=(\hbar/2)c_{i,\tau}^{\dagger}\boldsymbol{\sigma}_{\tau\tau^{\prime}}c_{i,\tau^{\prime}}$. The lattice operators $c_{n\tau}$ relate to $c_{{\boldsymbol{k}}\tau}$ in $\mathcal{H}_{\mathrm{NM}}$ via the Fourier transform $c_{n\tau}=(1/\sqrt{N_{N}})\sum_{{\boldsymbol{k}}}c_{{\boldsymbol{k}}\tau}e^{i{\boldsymbol{k}}\cdot{\boldsymbol{R}}_{n}}$, where $N_{N}$ is the number of NM unit cells and $\boldsymbol{k}$} runs over the 1BZ. Below, we consider a spatially uniform exchange coupling, i.e., $J_{\mathrm{sd},i}=J_{\mathrm{sd}}$, and assume that ${\boldsymbol{k}}\cdot{\boldsymbol{\iota}}_{i}\ll 1$, where ${\boldsymbol{\iota}}_{i}$ denotes the vector pointing from lattice site $i$ in the AFM to the center of the nearest Wannier orbital in the NM. We disregard Umklapp scattering between magnons and carriers.
@@ IV Experimental signatures
Similarly, the $z$-component of the spin current can be detected by interfacing the AFM with a magnetic metal and exploiting the anomalous inverse spin Hall effect (AISHE) [54, 55]. This effect is typically about an order of magnitude weaker than the conventional ISHE [56, 57, 58], which means that a ratio of $|I_{s,z}^{\pm}/I_{s,y}^{-}|\sim 10$ is required to obtain an AISHE voltage in the nanovolt range — achievable under sufficiently strong magnetic fields. For the two chiral states, one can expect a relative difference in the AISHE voltages on the order of $(I_{s,z}^{-}-I_{s,z}^{+})/I_{s,z}^{+}\sim 4\%$.
@@ V Summary and Conclusions
In conclusion, we have demonstrated that the ground-state vector spin chirality of a kagome AFM can be detected through measurements of the SSE. In the (-)-chiral state, the thermally pumped spin current develops a sizable in-plane component, which is absent in the (++)-chiral state. This component originates from an anisotropy-induced sublattice canting and is independent of the applied out-of-plane magnetic field. We also find a field-dependent out-of-plane spin current, whose magnitude differs between the two chiralities by about 4% due to their distinct magnon band structures. Together, the in-plane and out-of-plane SSE components provide clear signatures of the ground-state chirality, enabling real-time detection of chirality switching.

### v2

@@ Abstract
Non-collinear antiferromagnets (NCAFMs) are appealing for antiferromagnetic spintronics, as they combine the advantages of collinear antiferromagnets with novel emergent phenomena stemming from their complex spin structures. These phenomena are often associated with the intrinsic spin chirality, which characterizes the handedness of the ground-state spin configuration. Here, we investigate a kagome NCAFM interfaced with a normal metal and demonstrate that the ground-state vector spin chirality can be probed through measurements of the spin Seebeck effect (SSE). Starting from a microscopic spin Hamiltonian, we derive the corresponding bosonic Bogoliubov-de Gennes Hamiltonians for the two chiral configurations. Using linear response theory, we obtain a general expression for the spin current thermally pumped into the normal metal by the SSE. We show that a sizable in-plane spin current emerges exclusively in the negative-chiral state, providing a direct signature for real-time detection of chirality switching in kagome NCAFMs. In addition, we find a field-dependent out-of-plane spin current whose magnitude differs between the two chiralities by about 81%, reflecting their distinct magnon band structures.
@@ I Introduction
Chiral magnetic orders have garnered significant attention in the study of non-collinear antiferromagnets (NCAFMs) — antiferromagnetic systems in which the spin sublattices are ordered non-collinearly [5]. The complexity and chirality of their magnetic states are believed to give rise to unconventional magnetic, electronic, and optical phenomena, even in the absence of a net magnetization [6, 7, 8, 9, 10, 11, 12, 13]. A particularly attractive subclass of NCAFMs for exploring chirality-driven spin phenomena is the kagome antiferromagnets (AFMs). Kagome AFMs are two-dimensional (2D) spin systems composed of corner-sharing triangles and are considered among the most geometrically frustrated 2D spin lattices [14] (see Fig. 1). Important examples of materials that can be characterized as kagome AFMs are ultrathin films of the Weyl antiferromagnets [15] $\text{Mn}_{3}\text{X}$ (X = Ge, Sn, Ga, Ir) and iron jarosites [14], which consist of stacked kagome layers. In equilibrium, the three sublattice spins ${\boldsymbol{S}}_{1}$, ${\boldsymbol{S}}_{2}$, and ${\boldsymbol{S}}_{3}$ on each triangle are oriented at 120° to one another. The handedness of this spin configuration is described by the vector spin chirality ${\boldsymbol{K}}_{\mathrm{v}}={\boldsymbol{S}}_{2}\times{\boldsymbol{S}}_{1}+{\boldsymbol{S}}_{3}\times{\boldsymbol{S}}_{2}+{\boldsymbol{S}}_{1}\times{\boldsymbol{S}}_{3}$ [14]. For coplanar arrangements, ${\boldsymbol{K}}_{\mathrm{v}}$ points parallel or antiparallel to the z-axis, corresponding to the ground states of $(+)$-chirality and ()(-)-chirality, respectively (see Fig. 1). When subjected to an external magnetic field or an in-plane Dzyaloshinskii-Moriya interaction (DMI), the sublattice spins acquire a slight out-of-plane canting, leading to a non-coplanar spin texture. In this regime, the spin configuration also exhibits a finite scalar spin chirality expressed through the quantity $K_{\mathrm{s}}={\boldsymbol{S}}_{1}\cdot({\boldsymbol{S}}_{3}\times{\boldsymbol{S}}_{2})$ [14].
@@ II Theory
which includes the following contributions: an exchange interaction between nearest-neighbor spins ${\boldsymbol{S}}_{i}$ and ${\boldsymbol{S}}_{j}$ with coupling strength $J>0$, represented by $\mathcal{H}_{\text{E}}=J\sum_{\langle ij\rangle}{\boldsymbol{S}}_{i}\cdot{\boldsymbol{S}}_{j}$, easy-plane ($K_{\perp}>0$) and easy-axis ($K>0$) anisotropy energies, characterized by $\mathcal{H}_{\mathrm{A}}=\sum_{i}\left(K_{\perp}({\boldsymbol{S}}_{i}\cdot\hat{{\boldsymbol{z}}})^{2}-K({\boldsymbol{S}}_{i}\cdot\hat{{\boldsymbol{n}}}_{i})^{2}\right)$, an out-of-plane DMI between nearest neighbors in the form of $\mathcal{H}_{\mathrm{DM}}=\sum_{\langle i,j\rangle}D{\boldsymbol{\hat{z}}}\cdot({\boldsymbol{S}}_{i}\times{\boldsymbol{S}}_{j})$ with $\langle i,j\rangle\in\{\langle 1,3\rangle,\langle 2,1\rangle,\langle 3,2\rangle\}$ [36], and the Zeeman coupling to an external magnetic field $B$ applied perpendicular to the kagome lattice plane, $\mathcal{H}_{\mathrm{ext}}=-\sum_{i}\gamma B\hat{{\boldsymbol{z}}}\cdot{\boldsymbol{S}}_{i}$, where $\gamma$ is the gyromagnetic ratio. The in-plane easy-axes at the three spin sub-lattice sites are $\hat{{\boldsymbol{n}}}_{1}=[0,1,0]^{T}$, $\hat{{\boldsymbol{n}}}_{2}=[\frac{\sqrt{3}}{2},-\frac{1}{2},0]^{T}$, and $\hat{{\boldsymbol{n}}}_{3}=[-\frac{\sqrt{3}}{2},-\frac{1}{2},0]^{T}$, respectively. In what follows, it will often be convenient to represent a lattice site as $i\equiv(\kappa,\alpha)$, with $\kappa$ denoting the unit-cell index and $\alpha\in\{1,2,3\}$ the sublattice index.
@@
Figure 2: (color online). Magnon energy bands along high-symmetry lines in the first Brillouin zone for a. the $(+)$-chiral ground state and b. the ()(-)-chiral ground state. The inset diagrams highlight the band structures near the $\Gamma$ point along the $x$-axis. We have used the parameter values $\sqrt{2}a=3.785$ Å, spin $S=1$, $\gamma=1.76\cdot 10^{11}$ 1/Ts, $J=39.94~\mathrm{meV}/\hbar^{2}$, $K=5~\mathrm{meV}/\hbar^{2}$, $K_{\bot}=0$, $D=K/4\sqrt{3}$, and $B=5$ T.
Fig. 2 shows the magnon band structures calculated from Eq. (3) for the ($\pm$)-chiral states. The parameters correspond to $\text{Mn}_{3}\text{Ir}$ with lattice spacing $\sqrt{2}a=3.785$ Å [31], spin $S=1$, gyromagnetic ratio $\gamma=1.76\cdot 10^{11}$ 1/Ts, $J=39.94~\mathrm{meV}/\hbar^{2}$ [32, 31], $K_{\bot}=0$. $\text{Mn}_{3}\text{Ir}$ possesses three effective easy axes oriented approximately along the directions shown in Fig. 1 [31] with an associated effective anisotropy parameter of $K\approx 5~\mathrm{meV}/\hbar^{2}$ [32, 31]. For a $\text{Mn}_{3}\text{Ir}$ monolayer oriented along the kagome plane, these easy axes are expected to lie within the plane and coincide with those assumed in Eq. (1). For the out-of-plane DMI, we take the threshold value corresponding to the transition between positive and negative chiralities, $D=K/4\sqrt{3}$. Unless otherwise specified, these material parameters are used in all subsequent calculations. Additionally, for the band structures in Fig. 2, an external magnetic field of $B=5$ T is applied.
The main difference between the band structures of the $(\pm)$-chiral states arises near the $\Gamma$ point, as highlighted in the inset diagrams of Fig. 2. In the $(+)$-chiral state, the lowest band exhibits a finite energy gap at ${\boldsymbol{q}}=0$, whereas in the ()(-)-chiral state the lowest band approaches zero as ${\boldsymbol{q}}\to 0$. At ${\boldsymbol{q}}=0$, the band retains only a small finite value determined by the in-plane tilt angle $\alpha$, which vanishes in the strong exchange-coupling limit considered in Ref. [38, 39]. At finite temperature, this results in a larger magnon population in the lowest band of the ()(-)-chiral state compared to the $(+)$-chiral state, thereby leading to an amplification of the SSE, as we demonstrate below.
@@ III The Spin Seebeck Effect
with $J_{\mathrm{sd},i}\equiv\int{\rm d}{\boldsymbol{r}}J_{i}\rho_{i}({\boldsymbol{r}})|\Psi_{{\boldsymbol{R}}_{i}}({\boldsymbol{r}})|^{2}$ and ${\boldsymbol{s}}_{i}=(\hbar/2)c_{i,\tau}^{\dagger}\boldsymbol{\sigma}_{\tau\tau^{\prime}}c_{i,\tau^{\prime}}$. The lattice operators $c_{n\tau}$ relate to $c_{{\boldsymbol{k}}\tau}$ in $\mathcal{H}_{\mathrm{NM}}$ via the Fourier transform $c_{n\tau}=(1/\sqrt{N_{N}})\sum_{{\boldsymbol{k}}}c_{{\boldsymbol{k}}\tau}e^{i{\boldsymbol{k}}\cdot{\boldsymbol{R}}_{n}}$, where $N_{N}$ is the number of NM unit cells and $\boldsymbol{k}$} runs over the 1BZ. Below, we consider a spatially uniform exchange coupling, i.e., $J_{\mathrm{sd},i}=J_{\mathrm{sd}}$, and disregard Umklapp scattering between magnons and carriers. Furthermore, since the magnon wavelength is much larger than the lattice constant, we approximate $\exp(-i\Delta{\boldsymbol{k}}\cdot{\boldsymbol{\iota}}_{i})\approx 1$, where ${\boldsymbol{\iota}}_{i}$ is the vector pointing from lattice site $i$ in the AFM to the center of the nearest Wannier orbital $\Psi_{{\boldsymbol{R}}_{i}}({\boldsymbol{r}})$ in the NM and $\hbar\Delta{\boldsymbol{k}}$ denotes the typical momentum transfer experienced by electrons when scattering off a magnon. This approximation effectively amounts to neglecting the small spatial misalignment between the spin ${\boldsymbol{S}}_{i}$ and the center of the Wannier orbital $\Psi_{{\boldsymbol{R}}_{i}}({\boldsymbol{r}})$ in the Fourier transform of Eq. (6).
@@ IV Experimental signatures
Similarly, the $z$-component of the spin current can be detected by interfacing the AFM with a magnetic metal and exploiting the anomalous inverse spin Hall effect (AISHE) [57, 58]. This effect is typically about an order of magnitude weaker than the conventional ISHE [59, 60, 61], which means that a ratio of $|I_{s,z}^{\pm}/I_{s,y}^{-}|\sim 1$ is required to obtain an AISHE voltage in the nanovolt range — achievable under sufficiently strong magnetic fields. For the two chiral states, one can expect a relative difference in the AISHE voltages on the order of $(I_{s,z}^{-}-I_{s,z}^{+})/I_{s,z}^{+}\sim 81\%$.
@@ V Summary and Conclusions
In conclusion, we have demonstrated that the ground-state vector spin chirality of a kagome AFM can be detected through measurements of the SSE. In the (-)-chiral state, the thermally pumped spin current develops a sizable in-plane component, which is absent in the (++)-chiral state. This component originates from an anisotropy-induced sublattice canting and is independent of the applied out-of-plane magnetic field. We also find a field-dependent out-of-plane spin current, whose magnitude differs between the two chiralities by about 81% due to their distinct magnon band structures. Together, the in-plane and out-of-plane SSE components provide clear signatures of the ground-state chirality, enabling real-time detection of chirality switching.